OpenStudy (anonymous):

Evaluate 15 - |x| if x = -6. A. 21 B. -21 C. -9 D. 9

OpenStudy (hoblos):

15 - |-6| = 15-6 = 9

OpenStudy (anonymous):

the answer is 21 because when -6 is substituted as x in d bracket it will be 15-| -6| which will later result to 15 + 6=21 because - multiplied by - gives a +

OpenStudy (hoblos):

these are not brackets!!

OpenStudy (anonymous):

what re they?

OpenStudy (hoblos):

it is absolute value

OpenStudy (anonymous):

oh! ok

OpenStudy (anonymous):

thanks

OpenStudy (hoblos):

my pleasure XD